OpenStudy (anonymous):

Y= -16x^2+190+0 A= B= C= graph when done.

OpenStudy (anonymous):

I think you have to do the quadratic formula

OpenStudy (anonymous):

umm i forgot what that is.... the b= something right?

OpenStudy (anonymous):

in quadratic equation term with x^2 is A term with x is B term with constant is C

OpenStudy (anonymous):

yes -b + or - sqrt(b^2-4ac) all over at 2a

OpenStudy (anonymous):

okay thx

OpenStudy (anonymous):

in this case you can solve for your variables a b c

OpenStudy (anonymous):

\[b = -b \pm \frac{ \sqrt{b^2 - 4ac} }{ 2 }\]

OpenStudy (anonymous):

jazy its 2a

OpenStudy (anonymous):

@jazy

OpenStudy (anonymous):

I thought I put that. But yes, over 2a. (:

OpenStudy (anonymous):

\[b = -b \pm \frac{\sqrt{b^2 - 4ac}} { 2a }\]

OpenStudy (anonymous):

uughh this is sooo harrddd

OpenStudy (anonymous):

ax^2 + b^2 + c -16x^2 + 190 +0 what is a, b, and c?

OpenStudy (anonymous):

-16 190 0

OpenStudy (anonymous):

oh. i got -380

OpenStudy (anonymous):

how do you get -380

OpenStudy (anonymous):

you have the steps down, i think you know what you are doing

hartnn (hartnn):

also 'x' term is missing..

OpenStudy (anonymous):

oh snap i didnt do the bottom hold on

OpenStudy (anonymous):

so... id plug in the answer i got from the eqation as x?

OpenStudy (anonymous):

i got 0.95

OpenStudy (anonymous):

it right @jazy

OpenStudy (anonymous):

I think you had it right up to the -380. You have to divide by 2(-16) or -32. What is -380 /-32?

OpenStudy (anonymous):

you have it

OpenStudy (anonymous):

yea thats when i got 11.87

OpenStudy (anonymous):

then i pluged it in to the equation

OpenStudy (anonymous):

right(: what was the 0.95 for though?

OpenStudy (anonymous):

y= -16(11.87^2)............. the 11.87 is substituted for the x